# Multifractaility, topology and anomalous Hall conductivity on a 30 degrees twisted bilayer honeycomb lattice

### v2

@@ Abstract
We consider $30^{\circ}$ twisted bilayer formed by two copies of Haldane model and explore evolution of its properties with varying interlayer coupling strength. Specifically, we compute the system’s energy spectrum, its fractal dimensions, topological entanglement entropy, local Chern markers and anomalous Hall conductivity. We find that at weak interlayer coupling, the system remains gapped and retains topological properties of the isolated layers, but at strong interlayer coupling, the system forms a gapless multifractal state, which may also possess localized corner modes. We also establish that anomalous Hall conductivity can be used to characterize the system’s topological properties in the same way as a local Chern marker. We find that both the local Chern marker and anomalous Hall conductivity show sign of a phase transition, which is not visible from the energy spectrum.
@@ I Introduction
Motivated by this, we are interested whether it is possible to obtain a topological quasicrystal by smoothly deforming a crystal with known non-trivial topology. Specifically, we choose to consider 30 degrees twisted bilayer honeycomb lattice - a model which is currently actively studied in the context of twisted bilayer graphene [12, 13, 14, 15]. However we consider a twisted bilayer, in which each monolayer is described by gapped Haldane model [16, 17]. We explore our model’s evolution once the interlayer coupling is increased. We obtain that at weak coupling, the model retains all properties of Haldane model: bulk states remain extended and edge states still exists. But as the interlayer coupling increases, the model undergoes a transition into a strongly coupled state, which is gapless, and whose eigenstates are multifractal.
Characterizing topological properties of our model turns out to be challenging because most well-established topological concepts (e.g. Chern numbers) were developed in momentum space, i.e. only for crystalline systems. Nevertheless, we are still able to find out if the system is topological by computing its topological entanglement entropy [18]. Furthermore, we study its topological properties by computing ’local Chern marker’ [11]. Finally we numerically compute its anomalous Hall conductivity by using Kubo formula and find that it can be used to characterize topological properties of the quasicrystal in the same way as local Chern marker. We obtain that both the local Chern marker and anomalous Hall conductivity show the behavior of a phase transition at smaller $t_{inter}$ than the bulk energy gap closes.
This paper is organized as follows. In Sec. II we introduce our model, and in Sec. II.1 we describe its phase diagram, which consists of three phases: topological weakly-coupled, non-topological weakly coupled and strongly coupled multifractal. In Sec. II.2 we describe its topological properties. Specifically, in Sec. II.2.1 we present our results for its topological entanglement entropy and in Sec. II.2.2 we describe its topological properties using the ’local Chern markers’. In Sec. II.2.3 we discuss anomalous Hall conductivity. We summarize our findings in Sec. III .
@@ II Model
$H_{\alpha}$ $=$ $t_{intra}\sum\limits_{<ij>}c_{i,\alpha}^{\dagger}c_{j,\alpha}+t_{2}\sum\limits_{\ll i,j\gg}e^{-i\nu_{ij}\phi}c_{i,\alpha}^{\dagger}c_{j,\alpha}$
@@
$V_{1-2}$ $=$ $t_{inter}\sum\limits_{i,j}v_{i,j}c_{i,1}^{\dagger}c_{j,2}+h.c.$
[…]
$v_{i,j}=\left\{e^{-\frac{r_{ij}}{r_{0}}},\quad r<r_{max}\atop 0,\qquad r>r_{max}.\right.$
Here $r_{ij}$ is a distance between the sites $i,j$, and to simplify the model we neglect interlayer spacing. Also we introduced long-distance cutoff $r_{max}$ to decrease the complexity of our numerical calculations. We fix the numerical values of our parameters as $t_{intra}=1$, $t_{2}=1$, $\phi=\pi/2$, $r_{0}=1$, $r_{max}=2$ and explore its evolution over $m$, $t_{inter}$.
Figure 1: A schematic picture of the $30^{\circ}$ twisted bilayer of two honeycomb lattices. The top layer (blue) is formed by sites with coordinates $\vec{r}_{A,B}=\vec{a}_{1}n_{1}+\vec{a}_{2}n_{2}+\vec{c}_{A,B}$, where $\vec{a}_{1}=(\sqrt{3},0)$ , $\vec{a}_{2}=(\sqrt{3}/2,3/2)$, $\vec{c}_{A,B}=(0,\pm 1)$. The bottom layer (red) is obtained by rotating the top layer at 30 degrees. We consider an open system, whose sites are inside a square $-L<x,y<L$.
II.1 Phase diagram
To study the properties of our model, we compute its energy spectrum and wavefunctions numerically using exact diagonalization. Further, we characterize their spatial properties by computing fractal dimensions $D_{q}=\frac{\log(P_{q})}{(q-1)\log(N)}$, where $P_{q}$ is a participation ratio defined as $P_{q}=\sum_{i}(\psi^{\dagger}\psi)^{q}$, and $i=1\ldots N$ runs over all lattice sites.
Now let us look at the properties of our model once the interlayer coupling is turned on. We find that the phase diagram of the system in $m,t_{inter}$ space has three phases, which we label as ’topological weakly coupled’, ’non-topological weakly coupled’ and ’strongly coupled’ respectively (see Figs. 2). To be more specific, if we start from the topological phase of the Haldane model, we observe (see Figs. 3) that at small coupling, the bulk states are still separated by an energy gap and the edge states persist, but as $t_{inter}$ is increasing, the bulk energy gap shrinks. Once $t_{inter}$ reaches a critical value, the bulk gap closes, and the edge states disappear. As we increase $t_{inter}$ further, the system remains gapless. Similarly, if we start from the non-topological phase (see Fig. 3(b)), at small $t_{inter}$, the system remains gapped, and as $t_{inter}$ is increasing, the gap becomes smaller and eventually disappears. We also mention that before the gap closing, in both topological and non-topological phases, the system undergoes another phase transition, which does not manifest itself in the energy spectrum, but instead manifests only in the behavior of the local Chern marker and anomalous Hall conductivity. We discuss this in more details in Sec. II.2.2, II.2.3.
Figure 2: Schematic phase diagram of the $30^{\circ}$ degrees twisted bilayer Haldane model defined by the Eq. 1. The topological/non-topological gapped and gapless multifractal phases are obtained from the energy spectrum by exact diagonalization. The non-quantized gapped phases are distinguished by fluctuating local Chern marker and anomalous Hall conductivity.
@@ II.1 Phase diagram
Figure 3: (3(a)) Energies of states plotted against $m$ for various values of $t_{inter}$ and (3(b)) against $t_{inter}$ for values of $m$ in both topological and non-topological phases. We choose $L=30$, $q=2$. The color represents their fractal dimensions. At weak interlayer coupling, the spectrum is gapped for $m>1$, whereas for $m<1$ the gap is filled by the edge states, which have smaller fractal dimensions. At strong inerlayer coupling, the edge states disappear, and the spectrum becomes gapless.
To explore multifractal properties of our model, we plot average fractal dimension and its standard deviation for all states against the system size and a few values of $q$ (Fig. 4). At zero, as well as at small $t_{inter}$, we observe that as the system size increases, the average fractal dimension of all states in the system approaches $1$ and its standard deviation approaches 00. In other words, we confirm that in both topological and non-topological weakly coupled phases, the bulk eigenstates are fully extended in the same way as in a fully periodic system.
Figure 4: Average fractal dimensions $D_{ave}$ and their standard deviations $\Delta D$ over all states in a system plotted against its size $L$. We choose a list of $q=0.5,0.75,1.25,1.5,2.0,3.0,4.0$. In the weakly coupled phase, $D_{ave}$ converges to $1$, and $\Delta D$ converges to 00 - this means that all states are extended. In the strongly coupled phase, both $D_{ave}$ and $\Delta D$ converge to values between 00 and $1$ the system is multifractal.
Characterizing multifractal properties at large $t_{inter}$ turns out to be challenging because finite size corrections to fractal dimensions scale as $1/\log(L)$, and because of that it is impossible to obtain reliable results for any realistic system size, which can be handled numerically. Nevertheless, when we plot average fractal dimensions and their deviations as a function of the system size in the ’strongly coupled’ phase, we observe that the former converge to values smaller than $1$, and the latter converge to values greater than 00. Furthermore, we can see that as the system size grows, average fractal dimensions converge to different values for various $q$. Thus we confirm that eigenstates in the strongly coupled phase are multifractal - they are neither extended, nor localized.
To confirm that the system indeed undergoes a phase transition between the weakly and strongly coupled phases, we have to plot its derivatives of free energy and confirm that it diverges at the transition point. To be more specific, we consider a many body system at half filling and assume zero temperature. Simply speaking, we consider a many particle system, in which all states with negative energies are filled, but all states with positive energies are empty. Its free energy is just equal to the total energy, i.e. sum of energies of all filled states. First, we confirm that the topological phase transition is indeed a phase transition by plotting its second derivative as a function of $m$ (see Fig. 5(a)) and observing that it is discontinuous. Next, we plot the second derivative of free energy as a function of $t_{inter}$ (see Fig. 5(b)) and confirm that it has a peak at the transition point between the weakly and strongly coupled phases, whose magnitude grows with the system size.
@@
Figure 5: Second derivative of free energy as a function of $m$ (5(a)) and $t_{inter}$ (5(b)). The former is discontinuous at the phase transition points, and the latter is peaked. The colors refer to different lattice sizes: blue - $L=20$, yellow - $L=30$, green - $L=40$.
Finally let us briefly discuss the nature of the most localized states (i.e. states with the smallest fractal dimension) in the strongly coupled phase. We observe that such a phase on a square lattice hosts four corner modes, which were predicted in [16]. But even more interestingly, our system also holds other kinds of localized states. For instance, at $m=0$, we find that the states with the smallest fractal dimensions are localized in the center of the lattice (see Fig. 6). But at other values of $m$, there are states localized at specific locations in the bulk of the sample. We attribute the existence of such states due to the lack of translational symmetry. Indeed, in periodic crystals, isolated localized states cannot exist within the bulk of the material, but this does not have to be the case in a material which lacks translational symmetry. The same thing from a different perspective: in a periodic crystal, corner modes can be viewed as a consequence of bulk polarization [20], which is uniform in the bulk. However, bulk polarization does not have to be uniform in a system without translational symmetry. We leave the problem of building a more comprehensive theory of localized states in quasicrystals for future work.
((a)) $m=0.0$, $i=2236$
((b)) $m=0.0$, $i=4039$
((c)) $m=1.5$, $i=2762$
((d)) $m=1.5$, $i=2845$
Figure 6: $|\psi_{i}|^{2}$ for a few of the states among the ones with the smallest fractal dimensions. We assume $L=30$, $t_{inter}=5.0$. One can see that at $m=0$, there are states localized at the center of the lattice (6(a)) as well as the corner modes (6(b)). On the other hand, at $m=1.5$ we find states localized at other positions within the bulk of the lattice (6(c), 6(d)).
II.2 Topological properties
II.2.1 Topological entanglement entropy

### v3

# Localized states, topology and anomalous Hall conductivity on a 30 degrees twisted bilayer honeycomb lattice

@@ Abstract
We consider $30^{\circ}$ twisted bilayer formed by two copies of Haldane model and explore its evolution with varying interlayer coupling strength. Specifically, we compute the system’s energy spectrum, its fractal dimensions, topological entanglement entropy, local Chern markers and anomalous Hall conductivity. We find that at weak interlayer coupling, the system still has a bulk energy gap, topological edge states and retains topological properties of the isolated layers, but at strong interlayer coupling, this energy gap closes, but it is possible that another bulk gap may open. Such a system possesses multiple states localized at various locations of the lattice. We also establish that topological entanglement entropy and anomalous Hall conductivity can be used to characterize the system’s topological properties in the same way as a local Chern marker.
@@ I Introduction
Motivated by this, we are interested in studying a quasicrystal obtained by smoothly deforming a crystal with known non-trivial topology. Specifically, we choose to consider 30 degrees twisted bilayer honeycomb lattice - a model which is currently actively studied in the context of twisted bilayer graphene [12, 13, 14, 15]. However we consider a twisted bilayer, in which each monolayer is described by gapped Haldane model [16, 17, 18]. We explore our model’s evolution once the interlayer coupling is increased. We obtain that at weak coupling, the model retains all properties of Haldane model: its bulk spectrum remains gapped and the topological edge states persist. Thus we show that our system of two coupled copies of Chern insulators can be viewed as a topological quasicrystal. On the other hand, as the interlayer coupling becomes larger, the bulk energy gap closes, and the topological edge states disappear. Yet, since this system is non-periodic, its eigenstates are generally not extended. We obtain numerically that our bilayer possesses multiple kinds of states localized at various locations of the lattice. These include, but not limited to corner states, which were previously found in Refs. [17, 18]. We also observe that at large interlayer coupling, a new energy gap may open, but we argue that it is not topological.
In general, characterizing topological properties of a quasicrystal is challenging because most well-established topological concepts (e.g. Chern numbers) were developed in momentum space, i.e. only for crystalline systems. Nevertheless, we are still able to find out if the system is topological by computing its topological entanglement entropy [19] (we use computational methods previously considered in [20, 21, 22, 23]). Furthermore, we study its topological properties by computing ’local Chern marker’ [11]. Finally we numerically compute its anomalous Hall conductivity by using Kubo formula and find that it can be used to characterize topological properties of the quasicrystal in the same way as local Chern marker.
This paper is organized as follows. In Sec. II we introduce our model, and in Sec. II.1 we describe its phase diagram, which consists of multiple phases: topological weakly-coupled, non-topological weakly coupled and strongly coupled multifractal. The latter may have both gapped and gapless phases. We also comment on multifractal properties of our model (II.2). In Sec. II.3 we describe its topological properties. Specifically, in Sec. II.3.1 we present our results for its topological entanglement entropy and in Sec. II.3.2 we describe its topological properties using the ’local Chern markers’. In Sec. II.3.3 we discuss anomalous Hall conductivity. We summarize our findings in Sec. III .
@@ II Model
$H_{\alpha}$ $=$ $t_{intra}\sum\limits_{<ij>}c_{i,\alpha}^{\dagger}c_{j,\alpha}+t_{2}\sum\limits_{\ll i,j\gg}e^{-i\nu_{ij}\phi}c_{i,\alpha}^{\dagger}c_{j,\alpha}$ (2)
@@
$V_{1-2}$ $=$ $t_{inter}\sum\limits_{i,j}v_{i,j}c_{i,1}^{\dagger}c_{j,2}+h.c.$ (3)
Here $i,j$ numerate the lattice sites within each layer, <i,j><i,j> is a sum over the nearest neighbors, $\ll i,j\gg$ is a sum over the next-nearest neighbors. $t_{intra}$ is the nearest neghbors hopping (between A, B sites), $t_{2}$ is the next-nearest neighbors hopping (between A-A or B-B sites), $\nu=\pm 1$ depending on the relative direction between $i,j$, $\phi=\pi/2$, and $\epsilon=\pm 1$ for A(B) sites (see e.g. ref. [24] for more details). The interlayer potential has the form
$v_{i,j}=\left\{e^{-\frac{r_{ij}}{r_{0}}},\quad r<r_{max}\atop 0,\qquad r>r_{max}.\right$. (4)
Here $r_{ij}$ is a distance between the sites $i,j$, and to simplify the model we neglect interlayer spacing. Also we introduced long-distance cutoff $r_{max}$ to decrease the complexity of our numerical calculations. We fix the numerical values of our parameters as $t_{intra}=1$, $t_{2}=1$, $\phi=\pi/2$, $r_{0}=1$, $r_{max}=2$ (we checked that changing $t_{2}$, $r_{max}$ does not have qualitative impact on the model’s properties) and explore its evolution over $m$, $t_{inter}$. Since our 30 degrees twisted bilayer is a non-periodic system, we consider a system forming a square with equal lengths in both directions $L_{x}=L_{y}$ and fix open boundary conditions.
@@
((c))
Figure 1: A schematic picture of the $30^{\circ}$ twisted bilayer of two honeycomb lattices. Here we show its top-left corners with $L_{x}=20$ (1(a)), $L_{x}=30$ (1(b)), $L_{x}=40$ (1(c)). The top layer (blue) is formed by sites with coordinates $\vec{r}_{A,B}=\vec{a}_{1}n_{1}+\vec{a}_{2}n_{2}+\vec{c}_{A,B}$, where $\vec{a}_{1}=(\sqrt{3},0)$ , $\vec{a}_{2}=(\sqrt{3}/2,3/2)$, $\vec{c}_{A,B}=(0,\pm 1)$. The bottom layer (red) is obtained by rotating the top layer at 30 degrees. We consider an open system, whose sites are inside a square $L_{x}<x,y<L_{x}$. The cases (1(a), 1(b)) correspond to zigzag-armchair edges, and the case (1(c)) corresponds to the bearded-armchair edge.
II.1 Phase diagram
To study the properties of our model, we compute its energy spectrum and wavefunctions numerically using exact diagonalization. Further, we characterize their spatial properties by computing fractal dimensions
$D_{q}=\frac{\log(P_{q})}{(q-1)\log(N)},$ (5)
where $P_{q}$ is a participation ratio defined as $P_{q}=\sum_{i}(\psi^{\dagger}\psi)^{q}$, and $i=1\ldots N$ runs over all lattice sites. In a system of large spatial size, fractal dimension is expected to be close to $1$ for extended bulk states, $1/2$ for edge states and 0 for localized states (e.g. corner modes). In this section, we use $D_{q}$ to find the most localized states, which have the smallest $D_{q}$.
Once the interlayer coupling is turned on, the phase diagram of the system in $m,t_{inter}$ space has multiple phases, which we schematically outline on the Fig. 2 and which we obtained by running over different values of $m$ and $t_{inter}$ (see Figs. 3, 4). Specifically, if we start from the topological phase of the Haldane model, we observe (see Figs. 3) that at small coupling, the bulk states are still separated by an energy gap and the edge states persist, but as $t_{inter}$ is increasing, the bulk energy gap shrinks. Once $t_{inter}$ reaches a critical value ($t_{inter}\approx 1.5$ for $m=0$), the bulk gap closes, and the edge states disappear. As we increase $t_{inter}$ further, at small $m$, another energy gap appears (see Fig. 4(a), 4(b), 4(c) ). However, we do not observe it at larger $m$ (see Fig. 4(d) ). Similarly, if we start from the non-topological phase (see Fig. 4(f)), at small $t_{inter}$, the system remains gapped, and as $t_{inter}$ is increasing, the gap becomes smaller and eventually disappears. Near this phase transition, another gap may appear.
Let us look more closely at the properties of the states in our model. For simplicity, we focus on the case $m=0$. As we mentioned, at $t_{inter}=0$, our model is a superposition of two Haldane models, which in turn host bulk states separated by an energy gap and edge states within the bulk gap. However, we also found that our system hosts additional edge states within the bulk gap. We observe such states in the cases of both zigzag-armchair and bearded-armchair edges, but we find them to lie at slightly different energies. Specifically, in the case of zigzag-armchair edges, these edge states lie at $E\approx\pm 2.3$ (see Fig. 6(a)), whereas in the case of bearded-armchair edge, these states lie at energy $E\approx\pm 1.9$. One can see these states for various values of $L_{x}$ on the Fig. 5(a) - there bearded-armchair configuration corresponds to lattice sizes $7,10,13,16\ldots$, and zigzag-armchair is realized for all other integer values of $L_{x}$.
As we increase $t_{inter}$, the edge states in the bottom ‘band’ disappear. However, in the top ‘band’, they persist, and a few of them tend to become localized within the corners of our system (see Figs. 6(b), 6(e)). Specifically for $L_{x}=30$, we also find that a few of the topological edge states become localized at the corners (see Fig. 6(b); note that we could not establish connection between these modes and a shape of the lattice boundary). As we increase $t_{inter}$ further, the bulk gap closes, and the system becomes gapless. At even larger $t_{inter}$, another energy gap opens. Interestingly, we found that even though such a gap exists at any system size, its precise location and width varies a lot between different lattice sizes (see Figs. 5(c), 5(d)). We attribute it to the fact that since our system is quasiperiodic, the system becomes unique for each lattice size, and therefore it is extremely difficult to achieve ‘continuum limit’ (see also discussion in the Sec. II.2). We however, can see that as the system size increases, this gap becomes less pronounced (see Fig. 5).
As predicted in the previous works [16], our system hosts states localized at the corners. However, we argue that their location is not attributed to the bulk gap. Indeed, we can see that for $L_{x}=30$ (see Fig. 4(b)), the localized corner states are in the bulk gap for a range of $t_{inter}$ between $\approx 5$ and $10$, but at smaller or larger $t_{inter}$, they merge with the bulk spectrum. On the other hand, at $L_{x}=20$ (see Fig. 4(a)), we find that the corner modes lie entirely within the bulk spectrum away from the gap, but at the same time, within the bulk gap, there are extended states. We note that both $L_{x}=20,30$ in our example correspond to zigzag-armchair edges of the lattice, and even have a similar structure of the corners (see Fig. 1(a), 1(b) ). We observe qualitatively similar picture for $L_{x}=40$ with ‘armchair-bearded’ edge structure (Fig. 1(c)): localized corner modes lie within the bulk gap at $t_{inter}>5$, but within the bulk states at smaller $t_{inter}$. ((see Fig. 4(c))).
Even more interestingly, we observe numerically that our system hosts other kinds of localized states. For instance, even at large $t_{inter}$, our system still hosts states localized near edges (see Fig. 7(f) ). But most surprisingly, we find that, at $m=0$, the states with the smallest fractal dimensions are localized in the center of the lattice (see Fig. 7(a), 7(b)).
We also find that at $m\neq 0$, both corner and ‘center’ modes disappear, but instead our system hosts states localized at specific locations in the bulk of the sample (see Fig. 8). We attribute their existence to the lack of translational symmetry. Indeed, in periodic crystals, isolated localized states cannot exist within the bulk of the material, but this does not have to be the case in a material which lacks translational symmetry. The same thing from a different perspective: in a periodic crystal, corner modes can be viewed as a consequence of bulk polarization [25], which is uniform in the bulk. However, bulk polarization does not have to be uniform in a system without translational symmetry. We leave the problem of building a more comprehensive theory of localized states in quasicrystals for future work.
Figure 2: Schematic phase diagram of the $30^{\circ}$ degrees twisted bilayer Haldane model defined by the Eq. 1 obtained by exact diagonalization. At small $t_{inter}$, the model has exactly the same topological/non-topological phases as a single layer Haldane model. At larger $t_{inter}$, the bulk gap closes, and the in-gap edge states disappear. Simultaneously, another bulk gap may open.
((a))
@@ II.1 Phase diagram
Figure 3: (3(a)) Energies of states plotted against $m$ for various values of $t_{inter}$. We choose $L=30$, $q=2$. The color represents their fractal dimensions. At weak interlayer coupling, the spectrum is gapped for $m>1$, whereas for $m<1$ the gap is filled by the edge states, which have smaller fractal dimensions. At strong inerlayer coupling, the edge states disappear, and the spectrum becomes gapless.
@@
((c))
((d))
((e))
((f))
Figure 4: Energies of states plotted against $t_{inter}$ for $m=0$ and $L_{x}=20$ (4(a)), $L_{x}=30$ (4(a)), $L_{x}=30$ (4(c)) as well as $m=0.5$ (4(a)), $m=1.0$ (4(a)), $m=1.5$ (4(c)) at $L_{x}=30$. One can see that at small $t_{inter}$, the system retains properties of its counterpart at $t_{inter}=0$, but at large $t_{inter}$, the system either become gapless ($m=0.5,1$), or a new gap opens ($m=0,1.5$). Localized states have the smallest fractal dimensions and thus are shown in blue color.
((a))
((b))
((c))
((d))
Figure 5: Energies (vertical axis) and fractal dimensions (color) of the states plotted against the system size $L_{x}$ for $t_{inter}=0.0$ (5(a)), $t_{inter}=1.0$ (5(b)), $t_{inter}=5.0$ (5(c)) and $t_{inter}=10.0$ (5(d)). States with fractal dimensions smaller than $0.75$ are emphasized by larger dot size. One can see that at $t_{inter}=0.0$ and $t_{inter}=1.0$, topological edge states, as well as localized states within the bulk band exist for any $L_{x}$. In the strongly coupled phase ($t_{inter}=5,10$), one can see a new energy gap, but with increasing $L_{x}$ it becomes less pronounced. The most localized states (blue) may or may not lie within the energy gap.
((a))
((b))
((c))
((d))
((e))
((f))
Figure 6: Energies of states plotted against their fractal dimensions for $m=0$, $L_{x}=30$, $t_{inter}=0$ (6(a)), 11 (6(b)), 55 (6(c)) and $L_{x}=40$, $t_{inter}=0$ (6(d)), 11 (6(e)), 55 (6(f)). One can see that at $t_{inter}=0$ (6(a), 6(d)), the system forms a Chern insulator with topological edge states in the bulk gap, but in addition possesses edge states within the bulk band. As $t_{inter}$ increases (6(b), 6(e)), some of the edge states become localized near the corners. Such corner states persist at strong interlayer coupling (6(c), 6(f)). In addition, states localized at the center of the lattice emerge.
((a)) $i=1556$, E=3.4E=-3.4
((b)) $i=2236$, E=2.28E=-2.28
((c)) $i=1978$, E=2.65E=-2.65
((d)) $i=2075$, E=2.55E=-2.55
((e)) $i=4039$, $E=3.15$
((f)) $i=4045$, $E=3.52$
Figure 7: $|\psi_{i}|^{2}$ for a few of the localized states at $m=0$, $t_{inter}=5$, $L_{x}=30$. We observed states localized at the center of the lattice (7(a), 7(b)), forming a ring (7(c), 7(d)), a well as localized at the corners (7(e)) and at the edges (7(f))
((a)) $i=2760$, $E=1.3$
((b)) $i=2762$, $E=1.4$
((c)) $i=2764$, $E=1.4$
((d)) $i=2843$, $E=1.8$
((e)) $i=2845$, $E=1.85$
((f)) $i=2854$, $E=1.9$
Figure 8: $|\psi_{i}|^{2}$ for a few of the most localized localized states at $m=1.5$, $t_{inter}=5$, $L_{x}=30$. One can see that the states can be localized at various locations of the lattice.
II.2 Multifractal properties
To explore multifractal properties of our model, we plot average fractal dimension and its standard deviation for all states against the system size and a few values of $q$ (Fig. 9). At zero $t_{inter}$, we observe that as the system size increases, the average fractal dimension of all states in the system approaches $1$ and its standard deviation approaches 0. In other words, we confirm that in both topological and non-topological weakly coupled phases, the bulk eigenstates are fully extended as they are expected to be in a fully periodic system.
Figure 9: Average fractal dimensions $D_{ave}$ and their standard deviations $\Delta D$ over all states in a system plotted against its size $L$. We choose a list of $q=0.5,0.75,1.25,1.5,2.0,3.0,4.0$. In the weakly coupled phase, $D_{ave}$ converges to $1$, and $\Delta D$ converges to 0 - this means that all states are extended. In the strongly coupled phase, both $D_{ave}$ and $\Delta D$ converge to values between 0 and 11 - the system is multifractal.
Figure 10: Fractal dimensions $D$ of the states with the largest energies plotted against $1/\log(L_{x})$ for $q=0.5,0.75,1.25,1.5,2.0,3.0,4.0$. The wavefunctions were obtained using Lanczos diagonaliztion. We see that for $t_{inter}=0$, $D$ approaches $1$ with increasing $L_{x}$, but does not approach 11 for $t_{inter}\neq 0$.
Characterizing multifractal properties at non-zero $t_{inter}$ turns out to be challenging because finite size corrections to fractal dimensions scale as $\log(L) 11$1One can see this just by writing a wavefunctions as $\psi\sim C/\sqrt{L}$ and plugging into the Eq. 5, and because of that it is impossible to obtain reliable results for any realistic system size, which can be handled numerically. For this reason, we cannot say confidently whether average fractal dimensions approach $1$ at $t_{inter}=1$. However, when we plot average fractal dimensions and their deviations as a function of the system size in the ’strongly coupled’ phase ($t_{inter}=10$), we observe that the former converge to values smaller than $1$, and the latter converge to values greater than 0. Furthermore, we can see that as the system size grows, average fractal dimensions converge to different values for various $q$. Thus we can say fairly confidently that eigenstates in the strongly coupled phase are multifractal - they are neither extended, nor localized.
To complement our understanding of multifractaility in our system, we also study it using Lanczos diagonalization. This method makes it possible to consider much larger system sizes (up to $L_{x}\approx 800$). Specifically, for each $L_{x}$, we obtain eigenstates with the largest energies and compute their fractal dimensions (see Fig. 10). As previously, we can see that for $t_{inter}=0$, i.e. for periodic system, fractal dimension of the ‘top’ state approaches $1$ with increasing system size. On the other hand, for $t_{inter}\neq 0$, we can see that the fractal dimension of the ‘top’ state does not approach $1$. However, despite reaching huge lattice sizes, we are still not able to ensure that they are sufficient to make predictions about infinite-size properties. For instance, at $m=1.5$, $t_{inter}=10$, we observe that the wavefunction of the ‘top’ state remains rotationally-symmetric for sizes up to $L_{x}\approx 420$, whereas at larger sizes, starts getting localized near one of the corners (see Fig. 11). We believe that this is yet another manifestation of the fact that in a quasicrystalline system, continuum limit is reached extremely slowly, and even system with the largest sizes that can be handled by modern computers may still be far from it. In the future, it would be of interest to find means to study multifractality in our model analytically.
((a))
((b))
Figure 11: Wavefunctions of the states with highest energies obtained by Lanczos diagonalization for $m=1.5$, $t_{inter}=10$, $Lx=350$ (11(a)) and $L_{x}=450$ (11(b)).
II.3 Topological properties
II.3.1 Topological entanglement entropy